OpenStudy (tumblewolf):

Will fan and medal!! Linear Algebra help Write the standard form of the line that is parallel to 2x + 3y = 4 and passes through the point (1, -4). Include your work in your final answer.

OpenStudy (anonymous):

3y=4-2x

OpenStudy (anonymous):

Move the left side to the right to get standard form 2x+3y=4 -> 3y=4-2x remember to change the value to the opposite

OpenStudy (tumblewolf):

I've been playing around and keep getting 6x+3y=10 but that doesn't seem right to me

OpenStudy (anonymous):

doesnt seem like it, But i'm not sure how you would get it through 1 -4

OpenStudy (anonymous):

I just forgot how

OpenStudy (tumblewolf):

I did y-(-4)=-2/3(x-1) y=-2/3x+2/1 - 4 y=-2/3x-10/3 then for the conversion went y+2/3x=-10/3 2/3x+y = -10/3 6x+3y+-30/3 6x+3y=-10

OpenStudy (anonymous):

actually its very close

OpenStudy (tumblewolf):

Where did I go wrong?

OpenStudy (anonymous):

Hold on let me double check

OpenStudy (anonymous):

Yeah your right there

OpenStudy (anonymous):

What you wrote is a bit confusing tho

OpenStudy (anonymous):

I mean how you wrote it

OpenStudy (tumblewolf):

So i need to clarify how I wrote it? I have to submit the work along with the answer